If 9 is the solution of variable x in the equation ((5x - 7)/2) = y, then the value of y is 28.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement isFalse.
Consider the equation, ((5x - 7)/2) = y
Then, from the question it is given that the value of x = 9

So, (((5 × 9) - 7)/2) = y
((45 - 7)/2) = y
38/2 = y
y = 19

Thanks